OpenStudy (anonymous):

show me how to solve for x when 1/(x+1)^2 = [x/(x+1) - 2]/(x-1)

OpenStudy (anonymous):

i get stuck after multiplying both sides by x-1

OpenStudy (kinggeorge):

\[{1 \over{(x+1)^2}} ={{x \over x+1} - 2 \over x-1}\]Multiply both sides by \(x-1\) and \((x+1)^2\). You get\[x-1=(x+1)^2 \left({x \over x+1}-2\right)=x(x+1)-2(x+1)^2\]So you have \[x-1=x(x+1)-2(x+1)^2\]From here you should just be able to FOIL, combine like terms, and solve a quadratic equation for x.

OpenStudy (anonymous):

so x - 1 = -x^2 - 3x - 2 ... still confused bro sorry